OpenStudy (anonymous):

HELP PLEASE ?? I WILL MEDAL what is the value of d + e + f when d=20 , e= -4 and f= -2 ? A. -22 B. 18 C. 14 D. -26

OpenStudy (anonymous):

20 + (-4) + (-2) = 20 - 6 = 14 C. is the correct answer